OpenStudy (anonymous):

A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

I can't really help with this cuz I suck at it

OpenStudy (anonymous):

awww

Mehek (mehek14):

yea sowwy :(

OpenStudy (amistre64):

given 2 equations f and g newf = f + kg, for some constant k

OpenStudy (anonymous):

@amistre64 please explain that better :)

OpenStudy (amistre64):

you have 2 equations defined, call one of them f and the other g lets say g is a line: ax + by = c multiplying by some constant k gets us kg: kax + kby = kc solve each for y g: y = -ax/b + c/b kg : y = -kax/kb + kc/kb , all the ks cancel top to bottom so the kg = g is the same line

OpenStudy (anonymous):

Ok let me get this down :3

OpenStudy (amistre64):

how to explain the sum of the equations of a set is also an equation equal to the set.

OpenStudy (anonymous):

im still lost :( the way you word it makes me feel dumb

OpenStudy (amistre64):

lets do this, well just follow the instructions .... pick one of the equations

OpenStudy (anonymous):

umm ax+bx=c

OpenStudy (amistre64):

not quite what i meant, you posted 2 equations to play with, which pick one of them

OpenStudy (anonymous):

2x+8y=12

OpenStudy (amistre64):

now multiply it thru by some number, perferablly other than 0 or 1

OpenStudy (anonymous):

umm i'll multiply 8 by 2?

OpenStudy (amistre64):

lets say you want to use 2 as your multiplier 2(2x+8y = 12) --> 4x +16y = 24

OpenStudy (anonymous):

ohhh

OpenStudy (amistre64):

now add this new equation to the one you did NOT use. this lag is something else isnt it ...

OpenStudy (anonymous):

yes it is and 2x+2y(4x+16y) is that what you ment

OpenStudy (amistre64):

you gave me an equation you did not post you posted: 3x+8y=12 2x+2y=3 multiply one of them by some number, we did 2 2(3x+8y=12) --> 6x + 16y = 24 now add this to the other one (6x+16y)+(2x+2y) = (24)+(3)

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

27?

OpenStudy (amistre64):

in general spose we have 2 equations a1x + b1y = c1 a2x + b2y = c2 added up we have (a1+a2)x + (b1+b2)y = c1+c2 assume the solution set to the original 2 equations is (x1,y1) then it is also a soluton to the sum of the equations --------------------------- proof a1 x1 + b1 y1 = c1 a2 x1 + b2 y1 = c2 substitute them into equation 3 (a1+a2)x1 + (b1+b2)y1 = a1 x1 + b1 y1 + a2 x1 + b2 y1 (a1+a2)x1 + (b1+b2)y1 = (a1 +a2)x1 + (b1 +b2)y1 QED

OpenStudy (anonymous):

ok i do not understand that at all .-.

OpenStudy (amistre64):

its just a general proof that the solution to a system of equations is also a solution to any linear combination of the set of equations

OpenStudy (anonymous):

ahh so am is that it for this question?

OpenStudy (amistre64):

thtas the technical aspect of it yes, but the question is asking you to do something specific we obtained a new equation that is the sum of the other equations

OpenStudy (amistre64):

now we have to prove that the solution set to the first 2 is a solution to the one we created

OpenStudy (anonymous):

ok can we hurry this up ccuz at in 50 minutes i must leave (even though i don'tthink it will take this long)

OpenStudy (amistre64):

what is your solution set to the first 2 equations ?

OpenStudy (anonymous):

first solution for the first equation is 27??

OpenStudy (anonymous):

i havn't done the second equation

OpenStudy (amistre64):

we dont have time to fight the lag and try to cover how to find a solution to the first set of equations

OpenStudy (amistre64):

its assumed you know how to find the solution to the original setup to start with.

OpenStudy (amistre64):

if all 3 equations are equal with the same set of (x,y) values then they share the same solutions

OpenStudy (anonymous):

then their parallel lines

OpenStudy (amistre64):

good lcuk

OpenStudy (anonymous):

ok thanks